### v2

# CLEF 2024: Conference and Labs of the Evaluation Forum, September 09–12, 2024, Grenoble, France

Abstract
The rapid advancements in Large Language Models (LLMs) have opened new avenues for automating complex tasks in AI research. This paper investigates the efficacy of different LLMs-Mistral 7B, Llama-2, GPT-4-Turbo and GPT-4.o in extracting leaderboard information from empirical AI research articles. We explore three types of contextual inputs to the models: DocTAET (Document Title, Abstract, Experimental Setup, and Tabular Information), DocREC (Results, Experiments, and Conclusions), and DocFULL (entire document). Our comprehensive study evaluates the performance of these models in generating (Task, Dataset, Metric, Score) quadruples from research papers. The findings reveal significant insights into the strengths and limitations of each model and context type, providing valuable guidance for future AI research automation efforts.
Abstract
This paper explores the impact of context selection on the efficiency of Large Language Models (LLMs) in generating Artificial Intelligence (AI) research leaderboards, a task defined as the extraction of (Task, Dataset, Metric, Score) quadruples from scholarly articles. By framing this challenge as a text generation objective and employing instruction finetuning with the FLAN-T5 collection, we introduce a novel method that surpasses traditional Natural Language Inference (NLI) approaches in adapting to new developments without a predefined taxonomy. Through experimentation with three distinct context types of varying selectivity and length, our study demonstrates the importance of effective context selection in enhancing LLM accuracy and reducing hallucinations, providing a new pathway for the reliable and efficient generation of AI leaderboards. This contribution not only advances the state of the art in leaderboard generation but also sheds light on strategies to mitigate common challenges in LLM-based information extraction.
keywords:
[…]
1 Introduction
The unprecedented expansion of scientific publications [1, 2] in the field of artificial intelligence (AI) has created a substantial challenge in systematically tracking and assessing advancements. Leaderboards, which rank AI models based on their performance across various tasks and datasets, serve as a crucial mechanism for monitoring progress and driving competition. The extraction of state-of-the-art information, represented by (Task, Dataset, Metric, Score) quadruples, is vital for the maintenance and accuracy of these leaderboards.
The effectiveness of LLMs in information extraction (IE) tasks is deeply influenced by the context provided during the input phase. In this context, “context” refers to the segments of the scholarly articles from which the (Task, Dataset, Metric, Score) data is extracted. Unlike the concept of in-context learning, where models are primed with examples, our focus is on selecting relevant sections of text to guide the LLMs toward accurate and relevant information extraction while minimizing the risk of generating irrelevant or false information (hallucinations).
Figure 1: Main process overview
Existing research has shown that longer contextual inputs can lead to decreased model accuracy and higher rates of hallucination due to the potential distraction of irrelevant information [3, 4]. Therefore, identifying the optimal length and specificity of context is crucial [3]. This study explores three distinct types of contextual inputs to empirically evaluate their impact on the task of extracting leaderboards: DocTAET (Document Title, Abstract, Experimental Setup, and Tabular Information), DocREC (Results, Experiments, and Conclusions), and DocFULL (the entire document).
Our primary contributions are twofold. Firstly, we evaluate the performance of the latest LLMs, both open-sourced Llama-2 and Mistral 7B [5, 6] and proprietary GPT-4-Turbo and GPT-4.o [7], in generating structured summaries and classifying papers with and without leaderboard data. Our first research question (RQ1) encapsulates this: Which LLM provides the most accurate performance for generating structured summaries and for leaderboard/no-leaderboard classification?
Secondly, we focus on the precise extraction of individual elements, addressing our second research question (RQ2): Which LLM offers the best trade-off between precision and other performance metrics in few-shot and zero-shot settings? This question is critical, given the importance of precision in scholarly communications and the broader implications for the reliability of model outputs.
Additionally, we participate in the “SOTA? Tracking the State-of-the-Art in Scholarly Publications” shared Task 4 in the SimpleText [^1] https://simpletext-project.com/2024/en/ track of CLEF 2024 [8, 9]. The goal of the SOTA? shared task is to develop systems that, given the full text of an AI paper, can recognize whether the paper reports model scores on benchmark datasets and, if so, extract all pertinent (Task, Dataset, Metric, Score) tuples presented within the paper.
2 Related Work
The aspiration for automatically generated leaderboards to monitor advancements in scientific research has been a notable ambition within the scientific community. This initiative first gained traction through the analysis of citation networks, employing then-cutting-edge methodologies such as Markov Random Fields and others [10, 11].
The advent of transformer-based models [12] marked a significant leap forward, setting new benchmarks across a myriad of machine learning tasks. In this vein, researchers at IBM [13] utilized the Bert pre-trained model [14] within a Natural Language Inference (NLI) framework to discern entailment from the complete texts of research papers to their corresponding leaderboards.
In the realm of information extraction, the application of LLMs such as GPT-3.5 and LLAMA has demonstrated significant promise. A study by [15] highlighted the capabilities of GPT-3.5 and GPT-4 in clinical named entity recognition (NER), showcasing their adeptness at processing intricate clinical datasets with limited prerequisite training. Through strategic prompt engineering, these models exhibited remarkable improvements in performance for extracting medical entities from clinical documentation, reinforcing the potential of LLMs in executing complex NER tasks within the healthcare sector [15].
The emergence of LLMs, including ChatGPT, has sparked a reconsideration of specialized versus general-purpose training approaches in the context of LLMs. Building upon the foundation of utilizing LLMs for specialized information extraction tasks, [16] delves into the utilization of these models in the domain of virology. The study showcases how LLMs, specifically tuned for scientific content, can efficiently parse and extract virology-related information from a plethora of scientific publications. This research underlines the importance of fine-tuning and prompt engineering in enhancing the model’s ability to discern relevant scientific facts, contributing to the development of domain-specific leaderboards. The approach exemplified in this paper demonstrates an effective strategy for context selection in LLM-based leaderboard generation, emphasizing the necessity for domain-specific adjustments to maximize the accuracy and relevance of the extracted information.
In our previous study [17], we expanded on these findings by empirically investigating different ways to select context in creating leaderboards. We analyze the impact of tailored context cues and the integration of domain-specific knowledge on the precision and utility of automatically generated leaderboards. Our contributions included a thorough assessment of the latest LLMs, both open-source and proprietary, for assessing the effectiveness of different context selection strategies and the development of a novel methodology that significantly enhances the accuracy and relevance of LLM-based leaderboard generation. This study not only corroborates the pivotal role of context in leveraging LLMs for information extraction but also introduces innovative techniques that refine the process of generating insightful and reliable leaderboards in the scientific community.
Our Corpus
Train Test-Few-shot Test Zero-shot
Papers w/ leaderboards 7,987 753 241
Papers w/o leaderboards 4,401 648 548
Total TDM-triples 415,788 34,799 14,800
Distinct TDM-triples 11,998 1,917 1,267
Distinct Tasks 1,374 322 236
Distinct Datasets 4,816 947 647
Distinct Metrics 2,876 654 412
Avg. no. of TDM per paper 5.12 4.81 6.11
Avg. no. of TDMS per paper 6.95 5.81 7.86
Table 1: Our DocREC (Documents Result[s], Experimentation[s] and Conclusion) corpora statistics. The “papers w/o leaderboard” refers to papers that do not report leaderboard.
3 Methodology
This section details the methodology employed in our study, encompassing data collection, preprocessing, model selection, and evaluation metrics. The goal is to systematically evaluate the performance of four state-of-the-art LLMs—Mistral 7B, Llama-2, GPT-4-Turbo and GPT-4.o—across different context types for the task of leaderboard extraction.
3.1 Data Collection and Preprocessing
We utilized data that was previously published by [17] and re-released as the SOTA? shared task training corpus. The dataset consists of (T, D, M, S) annotations for thousands of AI articles available on PwC (CC BY-SA). These articles cover various AI domains such as Natural Language Processing, Computer Vision, Robotics, Graphs, Reasoning, etc., making them representative for empirical AI research. The specific PwC source was downloaded on December 09, 2023. The corpus comprised over 8,000 articles, with 7,987 used for training and 994 for testing, including 751 in the few-shot setting and 241 in the zero-shot setting. These articles, originally sourced from arXiv under CC-BY licenses, are available as latex source code, each accompanied by one or more (T, D, M, S) annotations from PwC. The articles’ metadata was directly obtained from the PwC data release, and the articles collection had to be reconstructed by downloading them from arXiv under CC-BY licenses.
After downloading the article’s source code (‘.tex‘), we needed to preprocess it to convert it to plain text. Sometimes, articles written in LaTeX are split into multiple files. To address this, we first created and executed a custom script to merge the project source code into a single LaTeX file corresponding to the arXiv ID of the paper. Next, we used another custom script to extract specific sections of the paper (DocTEAT or DocREC) from the `$arxiv\_id.tex`$ file, ensuring that the file remained compilable by LaTeX, which is necessary for our `$tex\_to\_text`$ parsing process.
[…]
Another important subset of our data, in addition to our base dataset reported in Table 1, was the "no leaderboards papers". We included a set of approximately 4,401 and 648 articles that do not report leaderboards into the train and test sets, respectively. These articles were randomly selected by leveraging the arxiv category feature, then filtering it to papers belonging to domains unrelated to AI/ML/Stats. These articles were annotated with the unanswerable label to finetune our language model in recognizing papers without (T,D,M,S) mentions in them.
We phrased the following question to formulate our task objective w.r.t. the (T, D, M, S) extraction target: What are the values for the following properties to construct a Leaderboard for the model introduced in this article: task, dataset, metric, and score? In essence, it encapsulates an IE task.
Instruction tuning [18, 19, 20, 21, 22, 23] boosts LLMs by providing specific finetuning instructions, improving adaptability and performance on new tasks [24, 25]. This method offers a more efficient approach than traditional unlabeled data methods [26, 27, 28, 29], allowing for versatile task prompting with single instructions.
In this vein, the “Flan 2022 Collection” [23] was a large-scale open-sourced collection of 62 prior publicly released datasets in the NLP community clustered as 12 task types, such as reading comprehension (RC), sentiment, natural language inference (NLI), struct to text, etc. It is the most comprehensive resource facilitating open-sourced LLM development as generic multi-task models. Importantly, and of relevance to this work, FLAN was not just a super-amalgamation of datasets encapsulating different learning objectives, but also included at least 10 human-curated natural instructions per dataset that described the task for that dataset. As such, we select a set of instructions to guide the LLM for our complex IE task from the FLAN collection. Specifically, we identified the applicable instructions to our task were those designed for the SQuAD_v2 [30, 31] and DROP [32] datasets. Specifically, 8 SQuAD and 7 DROP instructions were found suitable. The general characteristic of the selected instructions is that they encode a context and the SOTA task objective, and instruct the model to fulfill the objective. The context, in our case, is a selection from specific sections from the full-text of an article where (T, D, M, S) information is most likely shared. This is discussed next.
As introduced in our prior work [17], we compared the performances of the four state-of-the-art LLMs on the following three contexts:
3.2 DocTAET
Delineated in prior work [13], the context to the LLM comprises text selected from the (T)-title, (A)-abstract, (E)-experimental setup, and (T)-tabular information parts of the full-text. It yields an average context length of 493 words, ranging from a minimum of 26 words to a maximum of 7,361 words. These specific selections targeted the areas of the paper where the (T, D, M, S) were most likely to be found. An example of this context selection is illustrated in  Figure 2.
Figure 2: DocTAET representation of the paper title "Deformable Convolutions and LSTM-based Flexible Event Frame Fusion Network for Motion Deblurring" With Dashed lines representing Task, Dataset, Metrics, and Score present in the paper but not captured by paper with codes
3.3 DocREC
Introduced for the first time in this work, the DocREC context comprises text selected from the sections named (R)-results, (E)-experiments, and (C)-conclusions with allowances for variations in the three names. Complementary but still unique to DocTAET, the DocREC context representation aims to distill the essence of the research findings and conclusions into a succinct format. This context, ended by being much longer than DocTAET, yielded an average length of 1,586 words, with a minimum length of 27 words and a maximum length of 127,689 words. An example of this context selection is illustrated in  Figure 3.
Figure 3: DocREC representation of the paper title "Deformable Convolutions and LSTM-based Flexible Event Frame Fusion Network for Motion Deblurring". With Dashed lines representing Task, Dataset, Metrics, and Score present in the paper but not captured by paper with codes
3.4 DocFULL
As a last representation and to test the hypothesis that longer contexts which are not selectively tailored to suit the desired task objective tends to distract and thereby hinder the LLM performance, we used the full paper text as context. This approach entailed compiling the LaTeX source code of the document and translating its entirety into a plain text file. DocFULL ended producing the longest contexts compared to DocTAET and DocREC, in an average length of 5,948 words.
3.5 Models
In the period following the introduction of Flan-T5 [25], the community has witnessed the emergence of multiple advanced LLMs that have outperformed it. In this paper, we adopt the Flan-T5 fine-tuning paradigm and implement it in a comparative experimental setup on two of the latest state-of-the-art LLMs featured on the public LLM leaderboard.
In this work, we selected two open-sourced models as well as two closed-source models.
Mistral-7B. As the first open-sourced model, we selected Mistral-7B [6]. This model as the name suggested is a 7-billion-parameter language model optimized for performance and efficiency. It introduces the Grouped-Query Attention (GQA) for rapid inference and reduced memory requirements, and Sliding Window Attention (SWA) for handling long sequences with lower computational costs. The model surpasses existing models in benchmarks, including reasoning, mathematics, and code generation tasks. It also features fine-tuning capabilities for instruction following, achieving superior performance in human and automated benchmarks. Mistral 7B is designed for real-time applications, supports high throughput, and its architecture enables effective sequence generation with optimized memory usage. The model is released under the Apache 2.0 license, with its source code on Github, facilitating broad accessibility and application in various tasks.
Llama-2 7B. As the second open-sourced model, we selected the LLama-2 model [5]. The Llama-2 model is a collection of LLMs that range from 7 billion to 70 billion parameters, designed for both general and dialogue-specific applications. From the three available Llama-2 model checkpoints, i.e. 7B, 13B, and 70B, for comparability with our first select model, i.e. Mistral-7B, we choose the Llama-2 7B model. The Llama-2 models are fine-tuned for enhanced dialogue use cases and exhibit improved performance over existing open-source models in terms of helpfulness and safety, based on benchmarks and human evaluations. The Llama-2 family includes models optimized for different scales and introduces safety and fine-tuning methodologies to advance the responsible development of LLMs.
GPT-4-Turbo. As the first proprietary model, we leveraged GPT-4-Turbo, developed by OpenAI. GPT-4-Turbo is designed to offer enhanced performance and efficiency, providing faster response times and reduced computational costs. This model excels in various natural language processing tasks, including text generation, translation, summarization, and question answering. Optimized for high-throughput applications, GPT-4-Turbo maintains a high level of accuracy and consistency, making it suitable for real-time AI applications. Additionally, the model incorporates improvements in handling extensive contextual inputs, which enhances its capability in generating coherent and contextually relevant outputs.
GPT-4.o. For the proprietary model, we leveraged GPT-4.o, developed by OpenAI. GPT-4.o (“o” for “omni”) is designed to enable more natural human-computer interactions, accepting a wide range of input types including text, audio, image, and video, and generating outputs in text, audio, and image formats. The model is optimized for multimodal tasks, offering a significant advancement in vision and audio understanding compared to previous models.
We employed the QLORA (Quantum-enhanced Learning Optimization for Robust AI) [33] framework for fine-tuning our open-sourced models, leveraging its advanced optimization capabilities to enhance model performance. QLORA has been recognized for its innovative approach to integrating quantum computing principles with machine learning, offering a novel pathway to overcoming traditional optimization challenges.
4 Evaluations
Experimental setup. For training, we had one main experimental setting based on the 15 instructions. As elicited earlier in subsection 3.1, each of the 15 instruction were instantiated with the 7,987 (T, D, M, S) data instances and the SOTA question resulting in a total of 119,805 instances to instruction finetune out LLMs. In this scenario, we hypothesized that this repetition in the data instances across the instructions would cause the resulting model to overfit the training dataset. Thus to control for this, we applied the following experimental setup. Each instruction was instantiated with a random selection of only half the 7,987 (T, D, M, S) data instances resulting in a finetuning dataset of a sizeable 41,340 instances that had leaderboards. The papers w/o leaderboards were also similarly halved. As shown in Table 1, the few-shot test included 1,401 paper instances, comprising 648 papers without leaderboards and 735 papers with leaderboards. The zero-shot test set consisted of 789 papers, with 548 papers without leaderboards and 241 papers with leaderboards. Model hyperparameter details are in Appendix C. In terms of compute, all experiments including inference were run on an NVIDIA 3090 GPU. Training took  20 hours on the 50% sampled dataset, while inference lasted 10 minutes for  2k test instances.
Few-shot Zero-shot Model Rouge1 Rouge2 RougeL RougeLsum General -Accuracy  Rouge1 Rouge2 RougeL RougeLsum General -Accuracy Context Llama-2 7B 49.68 10.18 48.91 49.02 83.51 68.15 4.81 67.59 67.78 86.82 DocREC 49.70 17.62 48.81 48.81 83.62 62.75 10.88 62.07 62.18 86.22 DocTAET 5.38 0.79 4.96 5.13 57.54 7.55 0.71 7.24 7.35 37.80 DocFULL   Mistral 7B 55.46 14.11 54.54 54.64 88.44 72.98 6.87 72.42 72.35 92.40 DocREC 57.24 19.67 56.28 56.19 89.68 73.54 12.23 73.01 72.95 95.97 DocTAET 6.73 0.77 6.36 6.49 58.45 9.38 0.59 9.11 9.23 39.28 DocFULL   GPT-4-Turbo 52.64 5.82 51.99 51.76 60.89 72.80 2.66 72.35 72.09 77.06 DocREC 43.14 2.41 42.97 42.91 47.33 59.98 0.48 59.89 59.74 61.18 DocTAET 48.50 3.21 48.06 47.96 52.87 70.10 1.8 69.75 69.73 72.65 DocFULL   GPT-4.o 58.59 16.81 56.37 55.45 83.21 74.94 9.02 73.65 73.02 87.94 DocREC 52.10 13.72 50.77 49.26 80.63 69.59 8.81 68.57 67.43 87.56 DocTAET 55.41 17.82 53.01 51.79 79.56 70.05 10.89 68.42 67.51 78.95 DocFULL
Table 2: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.0 for the shared task, reported using the metrics proposed for the task. The output evaluations are conducted as a structured summary generation task (reported with ROUGE metrics) and as a binary classification task to distinguish between papers with and without leaderboards (reported as General Accuracy).
Metrics. We evaluated our models in two main settings. In the first setting, we applied standard summarization ROUGE metrics [34]. Furthermore, we also tested the models ability to identify papers with leaderboards and those without. This task was simple. For the papers with leaderboards, the model replied with a structured summary and for those it identified as without it replied as “unanswerable.” For these evaluations we applied simple accuracy measure. In the second setting, we evaluated the model JSON output in a fine-grained manner w.r.t. each of the inidividual (T, D, M, S) elements and overall for which we reported the results in terms of the standard F1 score and Precision score.
Few-shot Zero-shot Model Mode Task Dataset Metric Score Overall Task Dataset Metric Score Overall Context   Llama-2 7B Exact 20.93 13.06 13.96 3.04 12.75 13.97 6.83 11.72 2.61 8.78 DocREC Partial 31.37 22.50 21.99 3.46 19.83 24.05 16.6 18.28 3.10 15.51 Exact 29.53 16.68 20.02 1.14 16.84 21.75 11.26 16.99 0.77 12.69 DocTAET Partial 43.37 30.36 30.51 1.38 26.40 38.48 23.10 27.09 0.96 22.41 Exact 1.59 1.36 0.94 0.23 1.03 2.06 1.30 1.52 0.33 1.30 DocFULL Partial 2.29 1.82 1.68 0.37 1.54 3.36 2.49 2.49 0.54 2.22   Mistral 7B Exact 26.77 15.68 18.70 6.36 16.88 17.99 11.80 15.55 5.04 12.60 DocREC Partial 39.75 27.28 28.49 7.08 25.65 29.88 21.05 23.16 5.75 19.96 Exact 33.38 18.51 24.23 1.87 19.50 26.99 14.32 22.04 1.20 16.14 DocTAET Partial 46.35 32.75 34.16 2.25 28.88 44.90 27.29 32.23 1.41 26.46 Exact 0.81 0.57 0.57 0.56 0.63 0.22 0.33 0.33 0.76 0.42 DocFULL Partial 1.19 0.85 0.81 0.84 0.92 0.56 0.67 0.78 0.87 0.72   GPT-4-Turbo Exact 7.61 6.19 4.92 4.25 5.74 4.26 5.35 3.86 3.28 4.18 DocREC Partial 16.48 13.96 11.03 7.03 12.13 13.76 11.09 10.19 5.46 10.13 Exact 2.99 2.69 0.95 0.75 1.84 1.13 0.79 0.34 0.11 0.59 DocTAET Partial 6.22 5.42 3.03 1.63 4.08 2.72 1.59 1.59 0.11 1.5 Exact 3.38 3.16 1.98 2.48 2.75 2.45 2.98 1.81 2.77 2.5 DocFULL Partial 7.03 6.41 4.96 4.15 5.64 6.49 5.85 4.47 3.56 5.09   GPT-4.o Exact 16.14 16.11 15.50 10.76 14.63 16.04 15.05 17.43 10.38 14.72 DocREC Partial 38.40 32.63 29.35 15.20 28.90 37.23 31.16 29.97 14.96 28.33 Exact 14.10 12.76 9.91 2.11 9.72 13.78 10.25 11.01 2.36 9.35 DocTAET Partial 31.84 26.65 20.83 4.22 20.92 29.33 23.87 19.50 3.71 19.12 Exact 16.72 14.53 14.67 11.25 14.29 13.08 14.94 16.09 11.17 13.82 DocFULL Partial 36.56 31.0 27.61 16.50 27.93 35.59 28.28 27.38 14.80 26.52
Table 3: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.0 for the shared task, reported using the metrics proposed for the task. The evaluation considers the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output, reported in terms of F1 score.
Few-shot Zero-shot Model Mode Task Dataset Metric Score Overall Task Dataset Metric Score Overall Context   Llama-2 7B Exact 34.10 21.27 22.74 4.99 20.78 31.89 15.77 26.77 6.06 20.12 DocREC Partial 51.13 36.66 35.82 5.59 32.32 54.92 38.32 41.73 7.27 35.56 Exact 30.61 17.29 20.78 1.18 17.46 24.56 12.72 19.19 0.87 14.34 DocTAET Partial 44.96 31.48 31.66 1.43 27.38 43.46 26.09 30.60 1.09 25.31 Exact 34.69 29.59 20.41 5.10 22.45 32.20 20.34 23.73 5.08 20.34 DocFULL Partial 50.00 39.80 36.73 8.16 33.67 52.54 38.98 38.98 8.47 34.75   Mistral 7B Exact 37.65 22.15 26.38 8.94 23.78 35.68 23.40 31.02 9.98 25.02 DocREC Partial 55.90 38.52 40.18 9.95 36.14 59.25 41.73 46.20 11.46 39.66 Exact 39.48 21.89 28.66 2.21 23.06 38.46 20.41 31.41 1.71 23.00 DocTAET Partial 54.82 38.73 40.41 2.65 34.15 64.00 38.89 45.94 2.03 37.71 Exact 32.43 32.43 32.43 9.6 30.76 25.00 37.50 37.50 14.00 28.50 DocFULL Partial 71.43 48.65 45.95 14.52 45.13 62.50 75.00 87.50 21.62 61.66   GPT-4-Turbo Exact 30.46 24.80 19.67 17.30 23.06 20.57 25.96 18.66 15.94 20.28 DocREC Partial 65.96 55.95 44.07 23.89 47.47 66.51 53.85 49.28 21.67 47.83 Exact 32.79 29.51 10.27 8.20 20.19 29.41 20.59 8.82 2.94 15.44 DocTAET Partial 68.31 59.56 32.97 15.64 44.12 70.59 41.18 41.18 4.55 39.37 Exact 28.52 27.56 17.05 21.65 23.70 24.73 30.11 18.28 27.96 25.73 DocFULL Partial 59.32 55.91 42.64 30.69 47.14 65.59 59.14 45.16 30.91 50.20   GPT-4.o Exact 17.99 17.96 17.28 12.00 16.31 19.5 18.72 21.67 12.89 18.31 DocREC Partial 42.82 36.38 32.72 16.09 32.00 46.31 38.5 37.27 18.18 35.13 Exact 19.70 17.83 13.85 2.95 13.58 25.31 18.83 20.22 4.37 17.18 DocTAET Partial 44.50 37.25 29.11 7.78 29.66 53.86 43.83 35.80 9.07 35.64 Exact 19.18 16.67 16.84 12.90 16.40 15.30 17.48 18.83 12.05 16.17 DocFULL Partial 41.95 35.56 31.70 18.15 31.84 41.64 33.08 32.03 16.66 30.85
Table 4: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.o w.r.t. the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output in terms of Precision score.
Few-shot Zero-shot Model Mode Task Dataset Metric Score Overall Task Dataset Metric Score Overall Context   Llama-2 7B Exact 15.09 9.42 10.07 2.19 9.19 8.94 4.36 7.51 1.66 5.62 DocREC Partial 2.63 16.23 15.86 2.49 14.30 15.40 10.60 11.70 1.97 9.92 Exact 28.52 16.11 19.32 1.10 16.26 19.52 10.10 15.24 0.69 11.39 DocTAET Partial 41.89 29.33 29.43 1.33 25.50 34.53 20.73 24.31 0.85 20.10 Exact 0.81 0.69 0.48 0.12 0.53 1.06 0.67 0.78 0.17 0.67 DocFULL Partial 1.17 0.93 0.86 0.19 0.79 1.74 1.29 1.29 0.28 1.15   Mistral 7B Exact 20.77 12.14 14.49 4.93 13.08 12.03 7.89 10.38 3.37 8.42 DocREC Partial 30.84 21.12 22.07 5.49 19.88 19.98 14.07 15.45 3.84 13.33 Exact 28.91 16.03 20.99 1.62 16.89 20.79 11.03 16.97 0.93 12.43 DocTAET Partial 40.15 28.36 29.59 1.95 25.01 34.58 21.02 24.83 1.08 20.38 Exact 0.41 0.29 0.29 0.29 0.32 0.11 0.17 0.17 0.39 0.21 DocFULL Partial 0.6 0.43 0.41 0.43 0.47 0.28 0.34 0.39 0.44 0.36   GPT-4-Turbo Exact 4.35 3.53 2.81 2.42 3.28 2.37 2.98 2.15 1.83 2.33 DocREC Partial 9.42 7.98 6.30 4.12 6.95 7.67 6.18 5.68 3.12 5.66 Exact 1.56 1.41 0.50 0.39 0.96 0.58 0.40 0.17 0.06 0.30 DocTAET Partial 3.26 2.84 1.59 0.86 2.14 1.39 0.81 0.81 0.06 0.76 Exact 1.8 1.68 1.05 1.32 1.46 1.29 1.57 0.95 1.46 1.32 DocFULL Partial 3.73 3.40 2.63 2.23 3.00 3.42 3.08 2.35 1.89 2.68   GPT-4.o Exact 14.63 14.60 14.05 9.75 13.26 13.41 12.58 14.57 8.68 12.31 DocREC Partial 34.81 29.58 26.60 14.41 26.35 31.13 26.05 25.06 12.71 23.74 Exact 10.97 9.93 7.72 1.64 7.57 9.47 7.04 7.56 1.62 6.42 DocTAET Partial 24.79 20.75 16.21 2.89 16.16 20.15 16.40 13.39 2.33 13.07 Exact 14.82 12.88 13.00 9.97 12.66 11.42 13.05 14.05 9.76 12.07 DocFULL Partial 32.41 27.48 24.46 15.13 24.87 31.08 24.69 23.91 13.32 23.25
Table 5: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.o w.r.t. the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output in terms of Recall score.
4.1 Results and Discussion
This section analyzes the comparative performance of different contexts provided to LLMs, examining their impact on model precision and reliability across various tasks and settings. Additionally, we utilized a matching algorithm, fuzz.ratio, with a threshold of 50% for partial metrics to account for variations in how tasks, datasets, and metrics are reported in research papers compared to the pwc code dumps.
Results for RQ1: Performance in Generating Structured Summaries and Classification
  1. 1.
    Few-shot Performance (Table 2):
    • Llama-2 7B: Demonstrates consistent performance across different contexts (DocREC, DocTAET) with high General Accuracy in DocREC (83.51%) and DocTAET (83.62%). It shows substantial differences in ROUGE scores, especially in the ROUGE-2 and ROUGE-Lsum metrics, indicating its ability to capture more nuanced information.
    • Mistral 7B: Exhibits the highest General Accuracy in DocTAET (89.68%) and competitive ROUGE scores across all contexts. This model shows significant improvement over Llama-2 7B and GPT-4-Turbo in most metrics, showcasing the strengths of domain-specific fine-tuned open-source models.
    • GPT-4-Turbo: Shows moderate performance with lower General Accuracy in DocTAET (47.33%). However, it demonstrates notable improvements in ROUGE-L and ROUGE-Lsum, indicating its potential in capturing detailed summaries.
    • GPT-4.o: Shows strong performance in ROUGE metrics, especially in DocREC and DocTAET contexts, with high General Accuracy in DocREC (83.21%) and DocTAET (80.63%). However, the DocFULL context results of 79.56% which indicate the robustness of the DocREC context.
  2. 2.
    Zero-shot Performance (Table 2):
    • Llama-2 7B: Performs well with notable General Accuracy in DocREC (86.82%) and DocTAET (86.22%). However, there is a significant drop in performance in the DocFULL context (37.8% General Accuracy), likely due to the long and complex nature of the full document context which introduces distractions and challenges the model’s ability to maintain focus and coherence over extended text. This suggests that while the models perform well in more concise contexts, the added complexity and length in DocFULL can detract from their overall accuracy and effectiveness.
    • Mistral 7B: Outperforms other models in General Accuracy, especially in DocTAET (95.97%) and DocREC (92.40%). This indicates its robustness in zero-shot settings, making it a reliable choice for generating structured summaries without prior examples.
    • GPT-4-Turbo: Shows lower General Accuracy in zero-shot settings, particularly in DocREC (77.06%) and DocTAET (61.18%), highlighting the need for further optimization.
    • GPT-4.o: Continues to show competitive performance in DocREC (87.94%) and DocTAET (87.56%) contexts, reinforcing its capability in zero-shot scenarios.
Results for RQ2: Precision and Performance Trade-offs
  1. 1.
    Few-shot F1 Score analysis (Table 3):
    • Llama-2 7B: Shows a balanced performance between exact and partial metrics. For example, it achieves an Overall score of 26.40 in the partial mode under DocTAET context, indicating its ability to handle partial matches effectively.
    • Mistral 7B: Excels in both exact and partial metrics, with the highest Overall score of 28.88 in the partial mode under DocTAET context. This model provides the best trade-off between precision and other performance metrics.
    • GPT-4-Turbo: Shows moderate performance with an Overall score of 12.13 in the partial mode under DocREC context. It requires further optimization to compete with other models.
    • GPT-4.o: Demonstrates strong partial mode performance, with an Overall score of 28.90 under DocREC context.
  2. 2.
    Zero-shot F1 Score analysis (Table 3):
    • Llama-2 7B: Shows variability in performance with a notable Overall score of 22.41 in the partial mode under DocTAET context.
    • Mistral 7B: Continues to lead with the highest Overall score of 16.14 in the exact mode under DocTAET context, reinforcing its capability in zero-shot scenarios.
    • GPT-4-Turbo: Provides moderate performance in partial mode with an Overall score of 10.13 under DocFULL context, highlighting areas for improvement.
    • GPT-4.o: Provides solid performance in partial mode with an Overall score of 23.87 under DocTAET context.
  3. 3.
    Precision Scores (Table 4):
    • Llama-2 7B: Achieves moderate precision scores, particularly in the partial mode with an Overall score of 35.56 in zero-shot settings under DocREC context.
    • Mistral 7B: Dominates precision metrics with the highest Overall scores in both exact and partial modes across different contexts, demonstrating its precision and reliability.
    • GPT-4-Turbo: Shows competitive precision scores in zero-shot settings, particularly in the partial mode with an Overall score of 50.20 under DocFULL context.
    • GPT-4.o: Achieves moderate precision scores in few-shot settings, particularly in the exact mode, with an Overall score of 16.40 under DocFULL context.
  4. 4.
    Recall Scores (Table 5):
    • Llama-2 7B: Demonstrates notable recall scores in few-shot settings, particularly in the partial mode with an Overall score of 25.50 under DocTAET context. However, its performance varies significantly across different contexts and modes.
    • Mistral 7B: Leads in recall scores, achieving the highest Overall scores in both exact and partial modes across various contexts. Notably, it achieves an Overall score of 16.89 in the exact mode in few-shot setting and 12.43 in few-shot under DocTAET context.
    • GPT-4-Turbo: Shows moderate recall performance, with room for improvement, particularly in the zero-shot settings. Its best recall score is in the exact mode under DocREC context with an Overall score of 3.28.
    • GPT-4.o: Excels in recall scores, especially in few-shot settings, achieving an Overall score of 26.35 in the partial mode under DocREC context. It maintains competitive recall scores across different contexts and modes.
Score Extraction Analysis
When focusing on the extraction of score entities, GPT models, particularly GPT-4.o, demonstrate a clear advantage. As shown in Table 3, GPT-4.o outperforms other models in extracting scores with higher accuracy and precision. For example, in few-shot settings, GPT-4.o achieves an Overall score of 11.25 in the exact mode and 16.50 in the partial mode under the DocFULL context, significantly higher than Llama-2 7B and Mistral 7B. Similarly, in zero-shot settings, GPT-4.o maintains a competitive edge with an Overall score of 14.96 in the partial mode under DocREC context, surpassing the performance of the other models.
This superior performance in score extraction can be attributed to the advanced capabilities of GPT-4.o and its optimized architecture for handling diverse and long contexts. This makes it particularly effective in accurately recognizing and extracting numerical and textual data, which is crucial for generating accurate leaderboard metrics and scores.
Discussion
Open-source models like Mistral 7B demonstrate competitive and sometimes superior performance compared to proprietary models like GPT-4.o and GPT-4-Turbo. This is evident in both structured summary generation and classification tasks, particularly in the DocTAET context, where Mistral 7B consistently outperforms others.
The trade-off between precision and performance is well-balanced in DocREC on Mistral 7B, making it a reliable choice for applications requiring high precision. This is crucial for scholarly communications where accuracy and reliability are paramount.
Few-shot settings show that all models, including open-source ones, perform robustly, but Mistral 7B often leads in various metrics. This highlights the potential of fine-tuning open-source models to achieve high performance that surpasses closed-source models.
5 Conclusions
Our participation in the shared task has demonstrated that fine-tuning open-source models like Mistral 7B and Llama-2 7B can yield competitive, and in some cases superior, results compared to proprietary models such as GPT-4.o and GPT-4-Turbo. Throughout our experiments, the DocTAET context typically delivered dependable and accurate performance, while the DocREC context excelled in scenarios where precision is paramount. The interaction among different context types and their consequent impact on model performance provided valuable insights into ongoing research and practical deployment of LLMs.
The implications of our findings within the context of the shared task are substantial. They indicate that with meticulous context design and implementation, fine-tuned open-source LLMs are well-suited for tracking and synthesizing scientific progress. This capability enables the provision of current and nuanced leaderboards for any given academic field. The potential of this technology to support and augment the efforts of academics and policymakers is significant, heralding novel opportunities for automated or semi-automated leaderboard construction.
While we have made significant progress in elucidating and enhancing context selection for LLMs, several avenues for future exploration have emerged. Investigating hybrid context selection methods, domain-specific adaptations, and the integration of structured data could lead to even more sophisticated leaderboard generation. As LLMs continue to advance, optimizing their potential remains a dynamic and impactful field of research.
In conclusion, our involvement in the shared task has not only highlighted the effectiveness of fine-tuned open-source models but also emphasized the importance of strategic context selection in maximizing model performance. These insights contribute to the broader understanding of LLM capabilities and pave the way for future advancements in the field.
Acknowledgements
This work was jointly supported by the German BMBF project SCINEXT (01lS22070) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number: NFDI4DataScience (460234259).
Appendix A Instructions: Qualitative Examples
In this section, we elicit each of the instructions that were considered in this work as formulated in the FLAN 2022 Collection for the SQuAD_v2 and DROP datasets.
ID SQuAD_v2 Instructions DROP Instructions
1 Please answer a question about this article. If unanswerable, say "unanswerable". Answer based on context.
2 {Context} {Question} If unanswerable, say "unanswerable". Answer this question based on the article.
3 Try to answer this question if possible (otherwise reply "unanswerable"). {Context} {Question}
4 Please answer a question about this article, or say "unanswerable" if not possible. Answer this question: {Question}
5 If possible to answer this question, do so (else, reply "unanswerable"). Read this article and answer this question.
6 Answer this question, if possible (if impossible, reply "unanswerable"). Based on the above article, answer a question.
7 What is the answer? (If it cannot be answered, return "unanswerable"). Context: {Context} Question: {Question} Answer:
8 Now answer this question, if there is an answer (else, "unanswerable").
Table 6: Comparative Instructions for the SQuAD_v2 and DROP datasets.
Appendix B ROUGE Evaluation Metrics
The ROUGE metrics [34] are commonly used for evaluating the quality of text summarization systems. ROUGE-1 measures the overlap of unigram (single word) units between the generated summary and the reference summary. ROUGE-2 extends this to measure the overlap of bigram (two consecutive word) units. ROUGE-L calculates the longest common subsequence between the generated and reference summaries, which takes into account the order of words. ROUGE-LSum is an extension of ROUGE-L that considers multiple reference summaries by treating them as a single summary.
Appendix C Additional Data statistics and Hyperparameters
Our Corpus
Train Test-Few-shot Test Zero-shot
Papers w/ leaderboards 7,744 961 630
Papers w/o leaderboards 4,063 604 507
Total TDM-triples 612,709 74,475 76,936
Distinct TDM-triples 62,629 8,748 8,434
Distinct Tasks 1,365 432 339
Distinct Datasets 4,733 1,379 1,077
Distinct Metrics 2,845 850 825
Avg. no. of TDM per paper 5.1 5.1 6.7
Avg. no. of TDMS per paper 6.9 6.3 9.0
Table 7: Full Paper dataset statistics. The “papers w/o leaderboard” refers to papers that do not report leaderboard.
Our Corpus
Train Test-Few-shot Test Zero-shot
Papers w/ leaderboards 7,025 903 573
Papers w/o leaderboards 3,033 444 353
Total TDM-triples 515,203 63,041 61,474
Distinct TDM-triples 56,486 8,241 7,557
Distinct Tasks 1,232 417 323
Distinct Datasets 4,473 1,317 1,010
Distinct Metrics 2,687 812 778
Avg. no. of TDM per paper 5.0 5.0 6.7
Avg. no. of TDMS per paper 6.8 6.2 8.9
Table 8: DOCTEAT dataset statistics. The “papers w/o leaderboard” refers to papers that do not report leaderboard.
We used a context length of 2400 and based on GPU availability, a batch size of 2 and gradient_accumulation_steps of 4 were used, leading to a final batch size of 8. All experiments were run on five epochs and we used AdafactorSchedule and Adafactor optimizer [35] with scale_parameter=True, relative_step=True, warmup_init=True, lr=1e-4.